the question was use point-slope form to write the equation of the line with the gien properties. Write equation instandard form. it was "m = -3 passing through P(3,5)
Answer by mananth(16949)

"m = -3 passing through P(3,5)
y = mx+b ( plug the value of (x,y) to determine the intercept.
5=-3*3+b
5=-9+b
13=b
the equation will be y = -3x+13